### v4

# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful to measure the CME.
I Introduction
It has been predicted by quantum chromodynamics (QCD) that vacuum fluctuations can result in gluon fields of nonzero topological charges in local metastable domains. Interactions of quarks with such gluon fields can cause chirality imbalance, breaking the parity and charge-parity symmetries in those domains. Such chirality imbalance, under a strong magnetic field, would result in charge separation, a phenomenon called the chiral magnetic effect (CME) [1, 2, 3, 4].
Ultra-strong magnetic fields are presumably produced in non-central relativistic heavy ion collisions [5, 6]. The magnetic field is on average perpendicular to the reaction plane (RP, the plane span by the beam and the impact parameter direction of the collision). The CME-induced charge separation is along the direction of the magnetic field, and may be conveniently quantified by the $a_{1}$ variable in Fourier series of particle azimuthal distributions [7, 8],
$dN_{\pm}/d\tilde{\phi}\propto 1\pm 2a_{1}\sin\tilde{\phi}+2v_{2}\cos 2\tilde{\phi}$$\cdots\,$, (1)
where $\tilde{\phi}$ is the azimuthal angle of the particle momentum vector with respect to the RP and the subscript $\pm$’ indicates particle charge sign. The elliptic anisotropy $v_{2}$ harmonic is the leading modulation in particle distributions produced in relativistic heavy ion collisions [9]. Because of a vanishing mean $a_{1}$ due to random fluctuations, a commonly used observable is the three-point correlator [8],
$\gamma=\left\langle\cos(\phi_{\alpha}+\phi_{\beta}-2\psi)\right\rangle\,$, (2)
where $\phi_{\alpha}$ and $\phi_{\beta}$ are the azimuthal angles of two particles of interest (POIs), and $\psi$ is that of the RP (or participant plane because of fluctuations in the collision overlap geometry [10]). To cancel charge-independent backgrounds, such as effects from global momentum conservation, the difference between opposite-sign (OS) and same-sign (SS) correlators is used [11, 12],
$\Delta\gamma\equiv\gamma_{\rm os}-\gamma_{\rm ss}\,.$ (3)
The CME signal presented in the $\Delta\gamma$ observable would then be $2a_{1}^{2}$.
Definite signals of the CME have not yet been observed [11, 12, 13, 14, 15, 16]. The major difficulty is the large contamination in $\Delta\gamma$ from mundane QCD backgrounds. Those backgrounds arise from genuine two-particle correlations, such as correlations between daughter particles from a resonance decay, or between particles from the same jet or back-to-back dijet, coupled with elliptic flow anisotropies of those background sources [8, 17, 18, 19, 20, 21]. This background correlation can be schematically expressed as [8, 16, 22, 23]
$\Delta\gamma_{\rm res}=\left\langle\cos(\phi_{\alpha}+\phi_{\beta}-2\phi_{\rm res})\right\rangle v_{\rm 2,{\rm res}}\,$, (4)
where the subscript ‘res’ stands generically for correlated two-particle clusters such as resonances and jets, and $v_{\rm 2,{\rm res}}\equiv\left\langle\cos 2(\phi_{\rm res}-\psi)\right\rangle$ is the elliptic flow anisotropy of those background-contributing sources.
Large efforts have since been invested to eliminate or mitigate these backgrounds [24, 25, 26, 27, 28], including innovative observables [29, 30, 31, 32, 33, 34]. One of the techniques is the event-shape engineering (ESE) [35, 29, 30] method, grouping events into classes of different $v_{2}$ values relying on the dynamical fluctuations of $v_{2}$. A variation of this method is to select events according to the particle emission pattern, called the event-shape selection (ESS) method [36], relying on the statistical fluctuations of $v_{2}$. The main difference between the two methods is in the ways of quantifying the event shape. The ESE method [35] uses a variable calculated not from the POIs that are used for the physics measurement of $\Delta\gamma$ and $v_{2}$, while the ESS method uses a variable calculated from the POIs [36]. Each method analyzes $\Delta\gamma$ as a function of the measured $v_{2}$ in events binned according to the corresponding event-shape variable and projects $\Delta\gamma$ to $v_{2}=0$ to obtain the intercept, presumably more sensitive to CME signals than the overall $\Delta\gamma$ measurement and less sensitive (or ideally insensitive) to backgrounds.
The ESE method is relatively straightforward to comprehend. The ESS method is, however, complex because the POIs are used in both event selection and physics measurements. In this paper, we conduct a systematic study of the ESS and ESE methods for the CME search using several physics models as well as toy models of varying ingredients, with the goal to elucidate what the projected intercept entails from each method.
The focus of our study is placed on the effects of the event classifying variables on the intercept under various circumstances of the simulated events. We note, however, that the $v_{2}$ measured in experiments contains not only the flow anisotropy but also the so-called nonflow correlations, dependent of analysis methods. Nonflow correlations refer to those unrelated to the collision geometry common to the entire event, such as resonance decays and (di-)jet correlations [37, 38, 39, 40]. While we touch upon the nonflow issue in this article, a full and thorough discussion of nonflow effects on the ESE and ESS methods is outside the scope of the present study.
The rest of the paper is organized as follows. Section II describes the details of the ESE and ESS methods (and a related early method [24]) used for the CME search. Section III gives brief descriptions of the models used in this study. Section IV presents our model simulation results using ESS and ESE and discusses the findings. Finally, a summary is given in Section V. The appendix compiles all the $\Delta\gamma$ vs. $v_{2}$ plots from the simulations using the ESS and ESE methods.
II The Event-Shape Techniques
In ESE [35], events from a narrow centrality bin are grouped according to the ellipic flow vector magnitude $q_{2}$ calculated from particles (or other detected signals in experiment like energy depositions in calorimeters) that are different from the POIs. The event-by-event $q_{2}$ quantity is generally defined as
$q_{2}^{2}$ $=$ $\frac{1}{N}\left[\left(\sum_{i=1}^{N}\cos 2\phi_{i}\right)^{2}+\left(\sum_{i=1}^{N}\sin 2\phi_{i}\right)^{2}\right]$ (5)
$=$ $1+\frac{1}{N}\sum_{i\neq j}\cos 2(\phi_{i}-\phi_{j})\,$,
where $N$ is the number of particles in the event within a given kinematic region (we use “particles” in our description for convenience without loss of generality). One may use the normalized $\hat{q}_{2}$ quantity [36] for ESE,
$\hat{q}_{2}^{2}\equiv q_{2}^{2}/\left\langle q_{2}^{2}\right\rangle\,.$ (6)
The average of $q_{2}^{2}$ over all events, assuming Poisson fluctuations in multiplicity, is given by
$\left\langle q_{2}^{2}\right\rangle\approx 1+Nv_{2,q}^{2}\{2\}\,$, (7)
where
$v_{2}^{2}\{2\}=\left\langle\cos 2(\phi_{1}-\phi_{2})\right\rangle\,$, (8)
is the two-particle cumulant elliptic flow anisotropy ($\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the pair), and the subscript ‘$q$’ in $v_{2,q}$ indicates that it is calculated from the particles used for computing $q_{2}^{2}$.
In ESE, the $q_{2}^{2}$ is typically measured at forward and backward rapidities [29, 30], within a given narrow centrality range, and events are selected according to the measured value of $q_{2}^{2}$ to study $\Delta\gamma$ as a function of $v_{2}$ typically in the midrapidity region. The value of $q_{2}^{2}$ varies from event to event, and the variations are dominated by statistical fluctuations. These statistical fluctuations in $q_{2}^{2}$ at forward and backward rapidities are uncorrelated to those in $v_{2}$ (or any other quantities) at midrapidity, so the statistical fluctuation effect in $v_{2}$ is averaged to zero in those $q_{2}^{2}$-selected events. There are, however, also dynamical fluctuations affecting the value of $q_{2}^{2}$. The main source of dynamical fluctuations is the initial geometry, which can fluctuate event-by-event within a given (narrow) centrality bin [41, 42], or even at a centrality precisely determined by a particular centrality measure. Even if the colliding geometry was precisely fixed, there can still be fluctuations in the subsequent “hydrodynamic” evolution and interactions of the collision system. Because the sources of these dynamical fluctuations are common to the entire collision event, they affect $q_{2}^{2}$ and $v_{2}$ (and any other quantities) of the same event. Thus, the event classes selected by $q_{2}$ (or equivalently $q_{2}^{2}$ or $\hat{q}_{2}^{2}$) within a given narrow centrality range will have a different average $\left\langle v_{2}\right\rangle$ in the midrapidity region, or any other kinematic regions displaced from that for $q_{2}^{2}$ computation. In other words, because $q_{2}$ is dynamically correlated with $\left\langle v_{2}\right\rangle$, ESE is dividing events in groups of varying $\left\langle v_{2}\right\rangle$, which dynamically fluctuates from event to event. We note that since $\hat{q}_{2}^{2}$ and $\left\langle v_{2}\right\rangle$ are usually computed in separate $\eta$ regions, longitudinal decorrelation of flow [43, 44] can reduce the power of $\hat{q}_{2}^{2}$ in selecting dynamical fluctuations in $\left\langle v_{2}\right\rangle$. However, this effect is small because the flow decorrelation has been measured to be at most a few percent [45, 46].
Since the ESE method selects events according to dynamical fluctuations of the $\left\langle v_{2}\right\rangle$ of the POIs, decoupled from the $q_{2}^{2}$ computation, it is guaranteed that the average resonance $\left\langle v_{\rm 2,{\rm res}}\right\rangle$, or generally the elliptic anisotropies of all elements in the event including all CME background sources, are proportional to $\left\langle v_{2}\right\rangle$. Note that the final-state $\left\langle v_{2}\right\rangle$ contains contributions not only from primordial particles, but also from all decay products of resonances and clusters. Since the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $\left\langle v_{2}\right\rangle$ in those $q_{2}$-selected event classes are all connected to the initial geometry, they are all proportional to each other. This is illustrated in the cartoon in Fig. 1 (left), where the three ellipses illustrate event selections by $q_{2}$ and the hallow circles are the average $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ in each $q_{2}$-selected event class. In the ESE analyses for CME searches [29, 30], one analyzes $\Delta\gamma$ as a function of $\left\langle v_{2}\right\rangle$ in each event class and projects the $\Delta\gamma$ measurement to zero $\left\langle v_{2}\right\rangle$. The ESE method guarantees that the elliptic flows of the background sources also vanish at zero $\left\langle v_{2}\right\rangle$. In other words, a linear projection to $\left\langle v_{2}\right\rangle=0$ would be equivalent to a projection to $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$, where the CME backgrounds are zero. Since all those events are from a narrow centrality bin where the CME signal does not vary much, the intercept of the linear projection is presumably the CME signal. Note that there may still be complications from nonflow effects [47] depending on how $\Delta\gamma$ and $\left\langle v_{2}\right\rangle$ are measured, which is outside the scope of the present work.
Figure 1: Cartoons illustrating the various event-shape methods in terms of the event-by-event resonance $v_{\rm 2,{\rm res}}$ vs. final-state single particle $v_{2}$ in events within a narrow centrality bin. Here $v_{\rm 2,{\rm res}}$ and $v_{2}$ are taken to be the event-by-event elliptic shape variables, $\left\langle\cos 2(\phi-\psi)\right\rangle$, with respect to the event plane $\psi$ from another kinematic region different from the particles of interest (POIs). (Left panel) Event-Shape Engineering (ESE) where events are selected according to $q_{2}$ in kinematic region displaced from that of POIs’ (each $q_{2}$-class is indicated by the brown ellipse), in which the average $\left\langle v_{\rm 2,\rho}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$, unbiased, and the spread in $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ (indicated by the red circles) are due to dynamical fluctuations. (Center penal) Event-Shape (ES) method where events are binned in the observed elliptic shape variable $v_{2}^{\rm obs}$ of POIs (one bin is indicated by the brown rectangle), and the average $v_{\rm 2,{\rm res}}$ in those events (indicated by the red circles) are strongly biased, not strictly proportional to $v_{2}^{\rm obs}$ and with a positive $v_{\rm 2,{\rm res}}$ value at $v_{2}^{\rm obs}=0$. The wide range in $v_{2}^{\rm obs}$ (including negative values) is mainly due to statistical fluctuations. (Right panel) Event-Shape Selection (ESS) where events are selected according to $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs (each $\hat{q}_{2,{\rm pair}}^{2}$-class is indicated by the brown ellipse) in which the $v_{\rm 2,{\rm res}}$ is unnecessarily proportional to $v_{2}$ (indicated by the red circles) due to complicated biases from the $\hat{q}_{2,{\rm pair}}^{2}$ selection using the same POIs as for $v_{2}$. The wide spreads in $v_{2}$ and $v_{\rm 2,{\rm res}}$ are mainly due to statistical fluctuations inherited in the ESS method involving the same POIs. The green cross in each cartoon indicates the overall averages $\left\langle\left\langle v_{2}\right\rangle\right\rangle$ and $\left\langle\left\langle v_{\rm 2,{\rm res}}\right\rangle\right\rangle$ over the entire event sample.
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another kinematic region away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship between the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and the $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [48]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, but related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations, just like the $\hat{q}_{2}^{2}$ in the ESE method. However, unlike ESE, the $v_{2}$ variable in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events by ESS are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself and are calculated from the same POIs, so the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the momentum space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs. It is worthy to emphasize again that, while the event selection variable in both methods are dominated by statistical fluctuations event-by-event, these statistical fluctuations are canceled in $v_{2}$ in the ESE method because the calculations of $\hat{q}_{2}^{2}$ and $v_{2}$ use different sets of particles, but do not cancel (rather strongly correlated) in the $v_{2}$ in the ESS method because the calculations of $\hat{q}_{2,{\rm pair}}^{2}$ and $v_{2}$ use the same POIs.
Several combinations of $\hat{q}_{2,\textsc{poi}}^{2}$ and $v_{2}$ variables have been investigated in Ref. [36]. The $q_{2,\textsc{poi}}^{2}$ can be computed using single particle azimuthal angles of the POIs by Eq. 5. It can also be defined by pairs of POIs similar to Eq. 5; in this case, the $q_{2,{\rm pair}}^{2}$ of pairs is calculated by substituting the $i^{\rm th}$ particle’s $\phi_{i}$ by the azimuthal angle of the $i^{\rm th}$ pair in Eq. 5, and the sum runs over all pairs of POIs in the event. The azimuthal angle of a pair is defined to be that of the total momentum of the two POIs (vector sum of their momenta). The normalized $\hat{q}_{2,{\rm pair}}^{2}$ is then
$\hat{q}_{2,{\rm pair}}^{2}\equiv q_{2,{\rm pair}}^{2}/\left\langle q_{2,{\rm pair}}^{2}\right\rangle\,.$ (9)
where $\left\langle q_{2,{\rm pair}}^{2}\right\rangle\approx 1+N_{\rm pair}v_{2,{\rm pair}}^{2}\{2\}$. Similarly, one can define $v_{2}$ to be that of single particles of POIs or that of pairs of POIs; in the latter, one replace the azimuthal angle $\phi$ of a single particle by that of a particle pair in
$v_{2}=\left\langle\cos 2(\phi-\psi)\right\rangle\,.$ (10)
One can study the $\Delta\gamma$ as a function of $v_{2}$ of single POI particles (or $v_{2,{\rm pair}}$ of pairs of POI particles) selecting events on $\hat{q}_{2,\textsc{poi}}^{2}$ of single POIs (or $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs). It is found by the Anomalous-Viscous Fluid Dynamics (avfd) model study that, out of the four combinations, it is the best to use the combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ [36], because the linear intercept of $\Delta\gamma$ as a function of $v_{2,{\rm single}}$ of single POIs selecting events on $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs reflects most closely the true CME in avfd. However, since the same POIs are used for $\hat{q}_{2,{\rm pair}}^{2}$, $v_{2,{\rm single}}$ and $\Delta\gamma$, self-correlations are present in the measurement whose effects/biases are hard to discern. It is unclear how much background is remaining in the ESS intercept, and whether the remaining background is positive or negative. This is illustrated by the cartoon in Fig. 1 (right). In this work, we focus on the best combination of pair $\hat{q}_{2,{\rm pair}}^{2}$ and single $v_{2,{\rm single}}$ in ESS as found in [36], and examine what possible biases there may exist and the possible level of remaining background in the ESS intercept.
III Model setup and analysis details
We investigate four physics models that have been used to simulate heavy ion collisions. Three of them are dynamical models (avfd, ampt, epos4) and the fourth is a parameterization model (hydjet++). The avfd model is designed to study CME-related physics and has the capability to implement axial charge current and magnetic field. The other three models do not have those CME capabilities and are useful only for background studies, which is the main purpose of this work. The centralities of the models are determined by the charged hadron multiplicity within $|\eta|<0.5$ as in the STAR experiments [49].
For further insights, we also investigate background behaviors with two versions of a toy model. The inputs to the toy models are parameterizations of the measured data.
Figure 2: $m_{\rm inv}$ distributions. Shown are the per-event $m_{\rm inv}$ distribution of OS pairs subtracted by that of SS pairs from avfd with axial current densities $n_{5}/s=0,0.1$, and $0.2$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I simulation including primordial pions, and $K_{S}$ and $\rho^{0}$-resonance decay pions (lower center), and Toy Model II simulation including in addition decay pions from a mass continuum (lower right). Particles used to form the $m_{\rm inv}$ distributions are taken from all charged hadrons (defined to be $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in our study and treated as pions with pion mass) and are from the kinematic range of $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$. In the toy models all final-state particles are charged pions.
III.1 AVFD Model
The avfd model [50, 51, 52] is an anomalous fluid dynamics developed to describe the evolution of chiral fermion currents in the quark-gluon plasma (QGP) created in relativistic heavy ion collisions in addition to the normal VISHNU simulations [53]. The average magnetic field is calculated from the spectator protons, and it is nominally along the direction perpendicular to the RP. In the model, the magnetic field is quantified by event-by-event simulations [54] taking into account the fluctuations of the magnetic field direction with respect to the RP. A modest time evolution is assumed for the decreasing magnetic field with a typical lifetime comparable to the initial time of hydrodynamic evolution $\sim 0.6\text{ fm}/c$. The initial condition for the axial charge density ($n_{5}$) is dynamically generated in avfd to be proportional to the entropy density ($s$), and the strength is set via the proportionality coefficient $n_{5}/s$. Three $n_{5}/s$ values are used in the simulations, $n_{5}/s$=0, 0.1, and 0.2. The QGP medium evolution is simulated by viscous hydrodynamics to describe the bulk background in heavy ion collisions, with transport parameters for the diffusion coefficient as well as the relaxation time. A final hadronic stage after the hydrodynamic freeze-out is included with hadronic re-scatterings and resonance decays.
Au+Au collisions are simulated by avfd at $\sqrt{s_{{}_{NN}}}=200$ GeV in [34]. In this study, we focus on the centrality range of 30–40% as in [34]. Figure 2 (upper left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for the three values of $n_{5}/s$. The CME implemented in avfd affects the relative OS and SS pair $m_{\rm inv}$ distributions–the distribution flattens with larger $n_{5}/s$. This arises from contributions from back-to-back OS pairs (hence larger $m_{\rm inv}$) and near-side SS pairs (hence smaller $m_{\rm inv}$) as a result of the CME, resulting in enhancement at large $m_{\rm inv}$ and depletion at small $m_{\rm inv}$ for large $n_{5}/s$ compared to small one. Several resonance peaks are apparent in the $m_{\rm inv}$ distributions: $\rho^{0}$, $f_{0}(980)$, and $f_{2}(1270)$. No intrinsic mass width is implemented in avfd, so the resonances appear as $\delta$-functions. Weak-decay hadrons are treated as stable particles in avfd, so no $K_{S}$ peak is present in the $m_{\rm inv}$ distribution.
III.2 AMPT Model
ampt (A Multi-Phase Transport) is a parton transport model [55]. It consists of a fluctuating initial condition, parton elastic scatterings, quark coalescence for hadronization, and hadronic interactions. The initial condition of ampt is based on the hijing model [56]. The string melting version of ampt, which we use for our study, converts these initial hadrons into their valence quarks and antiquarks [57, 55]. The (anti-)quarks further evolve via two-body elastic scatterings, treated with Zhang’s parton cascade [58] with a total parton scattering cross section of 3 mb. After parton scatterings cease, a simple quark coalescence model is applied to convert partons into hadrons [55]. Subsequent interactions of those formed hadrons are modeled by a hadron cascade, including meson-meson, meson-baryon, and baryon-baryon elastic and inelastic scatterings [55]. The model has been extensively tested to reproduce the $p_{\rm T}$ spectra and flows of bulk particles.
Minimum bias (MB) Au+Au collisions are simulated by ampt at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. We terminate the hadronic interactions at a cutoff time of 30 fm/$c$ (by model default). Figure 2 (upper center) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $\rho^{0}$ peak is evident. No intrinsic mass width is implemented in ampt, so the resonance peak is a sharp $\delta$-function. Weak-decay hadrons are treated as stable particles in ampt, so no $K_{S}$ peak is present in the $m_{\rm inv}$ distribution.
III.3 EPOS4 Model
epos4 [59] is a Monte Carlo model to simulate high-energy proton-proton and nucleus-nucleus collisions, among others. At high energies many (nucleonic and partonic) interactions happen simultaneously at once, not sequentially. Parallel multiparton interactions are modeled via the pomeron exchange mechanism and result in complex configurations composed of many strings [60, 61]. The string decay products further interact, forming a medium treated by the core-corona picture followed by fluid dynamic evolution of the former and ordinary hadronic interactions in the latter [62]. The epos4 model has been extensively tested and tuned using proton-proton and heavy ion data.
MB Au+Au collisions are simulated by epos4 (version 4.0.0) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (upper right) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident. No intrinsic mass width is implemented in epos4, so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in epos4 and their decay pions are treated as “primordial” pions.
III.4 HYDJET++ Model
hydjet++ is an event generator to simulate heavy ion collisions by combining two independent components, namely, the soft physics part and the hard physics part [63, 64]. The former is determined by thermal equilibrium where hadrons are produced on the hypersurface represented by a parameterization of relativistic hydrodynamics with given freeze-out conditions. Chemical and kinetic freeze-outs are separate, and hadronic resonances are decayed. The hard physics part starts with an initial parton configuration from pythia [65], lets partons rescatter traversing the hot and dense nuclear medium and losing energy via collisions and radiative gluon emission, and follows with parton hadronization and particle formation. The model parameters are tuned to reproduce heavy ion data on charged particle multiplicity, $p_{\rm T}$-spectra and flow.
MB Au+Au collisions are generated by hydjet++ (version 2.4) at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. Figure 2 (lower left) shows the difference of the $m_{\rm inv}$ distributions between OS and SS pairs for 30–40% centrality, as an example. The $K_{S}$, $\rho^{0}$, and $f_{0}(980)$ resonance peaks are evident. No intrinsic mass width is implemented in hydjet++  so the resonance peaks are sharp $\delta$-functions. All $K_{S}$ particles are decayed in hydjet++ and their decay pions are treated as “primordial” pions.
III.5 Toy Model I
To gain insights, we have also generated particles using a toy model [48]. Toy models are useful and convenient for investigating the outcome of an analysis method given the known input to the model.
Two versions of the toy model are examined. In Toy Model I, we include primordial pions and two resonances, $K_{S}$ and $\rho^{0}$. The resonance mass is generated according to the Breit-Wignar distribution of proper mass and width [66]. We generate particles uniformly within the rapidity range of 1.5<y<1.5-1.5<y<1.5, taking event-by-event particle multiplicities according to Poisson statistics about the means of three times the measured average pseudorapidity densities [67, 49, 68]. The $p_{\rm T}$ spectra are parameterized based on the measured data [69, 70], namely,
$\frac{dN}{dp_{\rm T}}=\begin{cases}p_{\rm T}\left[\exp\left(\frac{m_{\rm T}}{T_{\rm BE}}\right)-1\right]^{-1}&\text{for pions, and}\\ p_{\rm T}\exp\left(-\frac{m_{\rm T}-m_{0}}{T}\right)&{\rm for\,resonances,}\end{cases}$ (11)
where transverse momentum $m_{\rm T}\equiv\sqrt{p_{\rm T}^{2}+m_{0}^{2}}$ with $m_{0}$ being the rest mass of the corresponding particle. The particle or resonance $p_{\rm T}$ is sampled from Eq. 11. At a given $p_{\rm T}$, the average $\left\langle v_{2}\right\rangle$ of pion, $K_{S}$, or $\rho^{0}$ meson is calculated based on the following parameterization of data motivated by the number-of-constituent-quark scaling [71],
$\left\langle v_{2}\right\rangle(p_{\rm T})=n_{q}\left[\frac{a}{1+\exp\left(-\frac{(m_{\rm T}-m_{0})/n_{q}-b}{c}\right)}-d\right]\,$, (12)
where $n_{q}$ is the number of constituent quarks for a given particle and, for our purpose, $n_{q}=2$. A 40% $v_{2}$ fluctuation is included about the parameterized $\left\langle v_{2}\right\rangle(p_{\rm T})$ event-by-event to obtain the final $v_{2}$ of the particle, based on which the azimuthal angle is generated about the fixed event plane at $\psi=0$.
The primordial pion, $K_{S}$, and $\rho^{0}$ input information are taken from experimental measurements of Au+Au collisions at 200 GeV in the centrality range of 30–40%, which we currently focus on. All parameters used in the toy model can be found in Ref. [48]. Figure 2 (lower center) shows the difference of the $m_{\rm inv}$ distributions between the OS and SS pairs from the generated toy-model data. We require that all resonances in our toy model (i.e. $K_{S}$ and $\rho^{0}$) to decay into a pair of charged pions. The decay is isotropic in the parent rest frame, and the decay pions are properly boosted to the lab frame. The $K_{S}$ and $\rho^{0}$ resonance peaks are evident.
As default only 2% of the measured $K_{S}$ abundance are generated in our toy model because this is the approximate fraction of $K_{S}$ whose decay daughters are both reconstructed as primordial pions in the STAR experiment. We vary this fraction to investigate its effect on our results.
III.6 Toy Model II
[…]
III.7 Analysis Details
Figure 3: $\hat{q}_{2}^{2}$ distributions. Shown are the single $\hat{q}_{2}^{2}$ and pair $\hat{q}_{2,{\rm pair}}^{2}$ distributions of POIs from avfd with $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). Shown in all panels are for the 30–40% centrality of Au+Au collisions, as examples. The $q_{2}^{2}$ (used in the ESE analysis) is calculated from particles in $|\eta|<0.3$, and the pair $q_{2,{\rm pair}}^{2}$ (used in the ESS analysis) is calculated from particles within $|\eta|<1$, both with $0.2<p_{\rm T}<2$ GeV/$c$.
POIs are assumed to be all charged hadrons, defined to be $\pi^{\pm}$, $K^{\pm}$, pp and $\bar{p}$ in our study, and are within the transverse momentum range of $0.2<p_{\rm T}<2$ GeV/$c$ and a certain central pseudorapidity ($\eta$) range, typical of midrapidity detectors such as the STAR experiment [72].
In the ESS method, the POIs are taken to be within $|\eta|<1$ (or $|\eta|<2$). The event selection quantity, pair $\hat{q}_{2,{\rm pair}}^{2}$, is computed by the two-particle cumulant method in Eqs. 5,9, always using all particle pairs of the same sample of POIs.
In the ESE method, the acceptance $|\eta|<1$ (or $|\eta|<2$) is divided into three subevents:
[…]
  • center subevent: $|\eta|<0.3$,
  • east subevent: $0.3<\eta<1$ (or $0.3<\eta<2$).
The POIs are taken from either the west subevent or the east subevent, but not from both. The event selection quantity $\hat{q}_{2}^{2}$ is computed by the two-particle cumulant method in Eqs. 5,6, using particles from the center subevent, $|\eta|<0.3$ (or from a forward/backward region $3<|\eta|<4$).
Figure 3 shows the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ and $\hat{q}_{2}^{2}$ distributions from the ESS and ESE methods, respectively. The two distributions are similar for each of the models; the $\hat{q}_{2}^{2}$ distribution ($|\eta|<0.3$) is broader than the $\hat{q}_{2,{\rm pair}}^{2}$ distribution ($|\eta|<1$), presumably due to larger fluctuations in the smaller acceptance. We have checked that the distribution of the single particle $\hat{q}_{2}^{2}$ calculated from $|\eta|<1$ is similar to the pair $\hat{q}_{2,{\rm pair}}^{2}$ from the same acceptance.
Figure 4: Average $v_{2}^{2}\{2\}$. Shown are the two-particle cumulant $v_{2}^{2}\{2\}$ of POIs and $v_{2,{\rm pair}}^{2}\{2\}$ of particle pairs of POIs as functions of centrality from avfd for $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). POIs are from acceptances $|\eta|<1$ (except for epos4, $|\eta|<2$) and $0.2<p_{\rm T}<2$ GeV/$c$. The large departure of $v_{2,{\rm pair}}^{2}\{2\}$ from $v_{2}^{2}\{2\}$, both calculated by the cumulant method (Eq. 8), is largely due to self-correlations where two pairs share a common particle. The issue is severe in peripheral collisions except for hydjet++ where the average multiplicity is significantly larger than the rest models; for reference, for 30–40% centrality of Au+Au, the average multiplicities within $|\eta|<0.5$ are 151, 169, 88, and 255 for avfd, ampt, epos4, and hydjet++, respectively.
The $\hat{q}_{2}^{2}$ variable calculated from single particles has been widely used. The $\hat{q}_{2,{\rm pair}}^{2}$ variable calculated from particle pairs is relatively new [36]. As seen from Eq. 7, the average $\left\langle q_{2}^{2}\right\rangle$ is related to the two-particle cumulant $v_{2}\{2\}$. To gain insights, we calculate by Eq. 8 the $v_{2}\{2\}$ of single POIs and the $v_{2,{\rm pair}}\{2\}$ of particle pairs of POIs; In the latter case, the $\phi_{1}$ and $\phi_{2}$ in Eq. 8 are the azimuthal angles of two pairs, each formed by two POIs. The $v_{2,{\rm pair}}\{2\}$ consists of a component from pairing of “random” (hydrodynamic) particle pairs or a hydrodynamic pair with a resonance-decay pair, and a component from two-resonance $v_{2}\{2\}$. The first component reflects, in a non-trivial way, the single particle $v_{2}$ [73]. The last component is simply resonance (background contributing source) $v_{2}\{2\}$. Both $v_{2}\{2\}$ and $v_{2,{\rm pair}}\{2\}$ contain nonflow effects, for example, from resonance decays (in the case of $v_{2,{\rm pair}}\{2\}$, the two pairs both contain daughters from the same resonance decay). Figure 4 shows $v_{2}\{2\}$ (from 2 POIs) and $v_{2,{\rm pair}}\{2\}$ (from 2 pairs of POIs) as functions of centrality in different simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200\text{ GeV}$. They roughly agree at high multiplicity (hydjet++, or central collisions of other models), and $v_{2,{\rm pair}}\{2\}$ becomes much larger than $v_{2}\{2\}$ at low multiplicity (peripheral collisions of models except for hydjet++). This discrepancy comes from the self-correlations where the two pairs share one common POI. To get a feeling of this effect, consider a simplified scenario with $\phi_{{\rm pair}}=(\phi_{1}+\phi_{2})/2$. For two pairs with indices $(1,3)$ and $(2,3)$ where particle 3 is shared, their correlation is $\cos 2[(\phi_{1}+\phi_{3})/2-(\phi_{2}+\phi_{3})/2]=\cos(\phi_{1}-\phi_{2})$, and the fraction of this case in the $v_{2,{\rm pair}}^{2}\{2\}$ calculation is the triplet multiplicity over the quadruplet one, which is $\sim 1/N$. Therefore, besides the common flow component, $v_{2,{\rm pair}}\{2\}$ has an extra part $\sim\langle\cos(\phi_{1}-\phi_{2})\rangle/N$, which accounts qualitatively for its excess over $v_{2}\{2\}$ when multiplicity is low.
In the ESS method, events are divided into bins of $\hat{q}_{2,{\rm pair}}^{2}$ of equal bin width 0.1. In the ESE method, events are grouped into five ranges in $q_{2}$, four equal size of width 0.5 from $q_{2}=0$ to 2, and the last range $q_{2}>2$. Note this is equivalent to grouping the events according to $\hat{q}_{2}^{2}$ with corresponding divisions.
For the ESS analysis, in each event class according to $\hat{q}_{2,{\rm pair}}^{2}$, we calculate the single particle $v_{2,{\rm single}}$ by Eq. 10 and the $\Delta\gamma$ correlator by Eqs. 2,3 using the EP method. The EP azimuthal angle is taken to be $\psi=0$ as fixed in models. The experimental analogy is data analysis using the first-order event plane reconstructed from spectator neutrons in zero-degree calorimeters (ZDC) [74].
Likewise, in each event class according to $q_{2}$ in the ESE analysis, we also use the known EP of $\psi=0$ in models to calculate $\left\langle v_{2}\right\rangle$ by Eq. 10 and $\Delta\gamma$ by Eqs. 2,3. Here, the POI is taken from the west or east subevent in calculating $\left\langle v_{2}\right\rangle$, and in calculating $\Delta\gamma$, both POIs are taken from either the west subevent or the east subevent. In addition, we use the two-particle cumulant (Eq. 8) to compute $v_{2}\{2\}$ where one particle is taken from the west subevent and the other from the east subevent, and the three-particle correlator [8] to compute
$\gamma=\left\langle\cos(\phi_{\alpha}+\phi_{\beta}-2\phi_{c})\right\rangle/v_{2}\{2\}\,$, (13)
where $\phi_{\alpha}$ and $\phi_{\beta}$ are the azimuthal angles of two POIs from the west (east) subevent and $\phi_{c}$ is that of a third particle from the other-side subevent, i.e. the east (west) subevent.
IV results and discussions
IV.1 Physics Models
Figure A.2 in the appendix shows the avfd simulation results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The $\hat{q}_{2,{\rm pair}}^{2}$ is computed from the same POIs used for the $\Delta\gamma$ and $v_{2,{\rm single}}$ measurements. Three values of $n_{5}/s$ are simulated. The results are consistent with those of Ref. [36]. In avfd and all other physics models we studied, $\Delta\gamma$ appears to be linear as a function of $v_{2,{\rm single}}$. We fit the measured range of $0<v_{2,{\rm single}}<0.2$ to obtain the intercept $\Delta\gamma_{\textsc{ess}}$ at $v_{2,{\rm single}}=0$. The intercept is sensitive to CME with the $v_{2}$-induced background significantly reduced. For $n_{5}/s=0$, the intercept is consistent with zero as one would expect if the background was completely removed. For $n_{5}/s=0.1$ and 0.2, the intercepts are found to be $1.34\times 10^{-4}$ and $5.70\times 10^{-4}$ with negligible statistical uncertainties, respectively. The CME signal is expected to scale with $(n_{5}/s)^{2}$. The ratio of the ESS intercept for $n_{5}/s=0.2$ to that for $n_{5}/s=0.1$ is $4.25\pm 0.13$, deviating from the expected ratio of 4 by about 2 standard deviations.
Figure A.2 shows the avfd results analyzed using the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$, separate from the POIs from $0.3<|\eta|<2$ (upper cut of 2 instead of typically 1 to increase statistics) that are used to calculate $\Delta\gamma$ and $v_{2}$. The range of dynamical variation of $v_{2}$ is significantly smaller than the range of variation in $v_{2,{\rm single}}$ of the ESS method, which is predominantly statistical. The intercept $\Delta\gamma_{\textsc{ese}}$ is obtained from a linear fit to all data points. The intercept for the $n_{5}/s=0$ case is consistent with zero, as expected. The intercepts for $n_{5}/s=0.1$ and 0.2 are $(1.63\pm 0.19)\times 10^{-4}$ and $(6.61\pm 0.19)\times 10^{-4}$, respectively. The statistical uncertainties are larger than those from ESS because of the larger range of projection to the intercept. The intercept ratio between $n_{5}/s=0.2$ and 0.1 is $4.06\pm 0.49$, consistent with the expected ratio of 4 although the uncertainty is large. It is interesting to note that the fit slopes from both ESS and ESE vary with $n_{5}/s$ significantly, whereas it is expected to measure the background strength. This is likely because the CME signal particles have altered the strength of background-contributing correlations in the final state [73] and thus affected the slope as well.
Figure A.3 shows the ampt results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are all negative; Negative intercepts were also found by the authors of Ref. [36, 75]. Since ampt does not have CME signals, the results suggest that the ESS intercept is sensitive not only to CME signals but also to backgrounds. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV; this is expected because the $\Delta\gamma$ background is approximately inversely proportional to multiplicity.
Figure A.5 shows the ampt results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ calculated from particles within $|\eta|<0.3$ whereas POIs are taken from $0.3<|\eta|<1$. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions. We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities $3<|\eta|<5$ and $2<|\eta|<3$ for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.5 and are found to be similar to those with midrapidity $\hat{q}_{2}^{2}$ but with larger uncertainties.
Figure A.7 shows the epos4 results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The POI and $\hat{q}_{2,{\rm pair}}^{2}$ pseudorapidity acceptance is taken to be $|\eta|<2$ to increase statistics. The intercepts appear to be positive, opposite to those in ampt. This again suggests that the ESS intercept may be sensitive to backgrounds since epos4 does not have CME signals. There is an indication of an increasing fit slope with decreasing centrality; however, the uncertainties are large.
Figure A.7 shows the epos4 results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from $|\eta|<0.3$. The POI pseudorapidity acceptance is taken to be $0.3<|\eta|<2$ to increase statistics. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions.
Figure A.9 shows the hydjet++ results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are consistent with zero at 200 GeV but are all positive at 27 GeV. Since hydjet++ does not have CME signals, the results again indicate that the ESS intercept is sensitive not only to CME signals but also to backgrounds. The hydjet++ results further suggest that the degree of the sensitivity to backgrounds depends on the collision energy, or details of the event content. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV, as expected, similar to the ampt results.
Figure A.9 shows the hydjet++ results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The intercepts are all consistent with zero, as expected, except the 20–30% and 40–50% centralities at 27 GeV where the intercept is displaced from zero but only with less than 2 standard deviations. The fit slope increases with decreasing centrality.
We have also used the three-particle correlator method to calculate $\Delta\gamma$ and the two-particle cumulant to calculate $v_{2}\{2\}$, without relying on the known EP at $\psi=0$. The results are thus subject to nonflow effects. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The results are consistent with those in Fig. A.9 calculated with the known $\psi=0$, with somewhat larger statistical uncertainties. This suggests that the nonflow effects are relatively small in hydjet++ within the statistical uncertainties of the current study.
We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities of $3<|\eta|<5$ and $2<|\eta|<3$ for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The intercepts are all consistent with zero, except the 40–50% centrality bin at 27 GeV where the intercept is positive but within 2 standard deviations from zero.
IV.2 Discussions on Physics Model Results
Figure 5: ESE intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESE method for various physics models. For avfd ($n_{5}/s=0$), the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are averaged over the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV ($\left\langle\Delta\gamma_{\textsc{ese}}/\Delta\gamma\right\rangle$ does not necessarily equal $\left\langle\Delta\gamma_{\textsc{ese}}\right\rangle/\left\langle\Delta\gamma\right\rangle$). The POI $\eta$ acceptance is $0.3<|\eta|<1$ for avfd, ampt, and hydjet++, and $0.3<|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $0.2<p_{\rm T}<2$ GeV/$c$ for all models. The $q_{2}^{2}$ is calculated using particles from $0.2<p_{\rm T}<2$ GeV/$c$ and $|\eta|<0.3$ for all models (and also forward/backward rapidities for ampt and hydjet++: $3<|\eta|<5$ at 200 GeV and $2<|\eta|<3$ at 27 GeV). Several points are missing as they are outside the frame but consistent with zero with large uncertainties.
Figure 6: ESS intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESS method for various physics models. For avfd ($n_{5}/s=0)$, the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are averaged over the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV ($\left\langle\Delta\gamma_{\textsc{ess}}/\Delta\gamma\right\rangle$ does not necessarily equal $\left\langle\Delta\gamma_{\textsc{ess}}\right\rangle/\left\langle\Delta\gamma\right\rangle$). The POI $\eta$ acceptance is $|\eta|<1$ for avfd, ampt, and hydjet++, and $|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $0.2<p_{\rm T}<2$ GeV/$c$ for all models. The $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs.
We summarize the intercepts $\left\langle\Delta\gamma_{\textsc{ese}}\right\rangle$ and the intercepts divided by the overall $\Delta\gamma$ magnitude $\left\langle\Delta\gamma_{\textsc{ese}}/\Delta\gamma\right\rangle$ averaged over the 20–50% centrality range from the physics model studies by the ESE method in Fig. 5. The ESE intercepts are mostly limited by large statistical uncertainties. The large statistical uncertainties arise from the narrow dynamical range of $\left\langle v_{2}\right\rangle$ based on the $\hat{q}_{2}$ event selection in a given narrow centrality bin (see, particularly, Fig. A.5 lower panels, Fig. A.5, and Fig. A.7) and from the long extrapolation lever arm to the $\left\langle v_{2}\right\rangle=0$ intercept. This is particularly true for 27 GeV where some of the data points fall outside the graph frame of Fig. 5. The avfd, ampt, and epos4 models are computationally expensive, and it is difficult to accumulate statistics for a qualitative improvement beyond the present precision.
As discussed in Sect. II, the flow-induced background must be zero at the ESE intercept $\Delta\gamma_{\textsc{ese}}$ in the absence of nonflow. In most of our model studies and for all the model studies presented in Fig. 5, we used the known reaction plane $\psi=0$ in computing the $v_{2}$ and $\Delta\gamma$ variables, so nonflow is absent from these observables. While the statistical uncertainties are large, the ampt and hydjet++ $\Delta\gamma_{\textsc{ese}}$ using $\hat{q}_{2}^{2}$ from the $|\eta|<0.3$ range seem to be nonzero. The reason is likely due to nonflow effects caused by the ESE event selection using the midrapidity $\hat{q}_{2}^{2}$: because the $\eta$ distance is small, there can be nonflow correlation between a POI and a particle used in the $\hat{q}_{2}^{2}$ calculation. Such a correlation could bias the $\left\langle v_{2,{\rm res}}\right\rangle$ values of resonances contributing to CME backgrounds to be not strictly proportional to the final-state particle $\left\langle v_{2}\right\rangle$, resulting in a nonzero $\Delta\gamma_{\textsc{ese}}$ intercept. To reduce this effect, we also calculate $\hat{q}_{2}^{2}$ using particles at forward/backward pseudorapidities in ampt and hydjet++, the results of which are shown by the red points in Fig. 5. These intercepts $\Delta\gamma_{\textsc{ese}}$ are consistent with zero, confirming our hypothesis, though the uncertainty is large for the ampt result at 27 GeV.
Figure 6 summarizes the physics model results from the ESS method on the intercepts $\left\langle\Delta\gamma_{\textsc{ess}}\right\rangle$ and the intercepts divided by the overall $\Delta\gamma$ magnitude $\left\langle\Delta\gamma_{\textsc{ess}}/\Delta\gamma\right\rangle$, averaged over the 20–50% centrality range. As discussed in Sect. II, the ESS method uses the POIs to compute the event selection $\hat{q}_{2,{\rm pair}}^{2}$ variable, which are automatically (self-)correlated with the $\Delta\gamma$ and $v_{2,{\rm single}}$ of POIs. Thus, the $\Delta\gamma$ and $v_{2,{\rm single}}$ are affected by statistical fluctuations in $\hat{q}_{2,{\rm pair}}^{2}$ that is used to select events. The wide range of $v_{2,{\rm single}}$ and $\Delta\gamma$ are a result of these statistical fluctuations by insisting on certain values of the statistically fluctuating $\hat{q}_{2,{\rm pair}}^{2}$. The $v_{2,{\rm single}}$ value can go to zero as shown in Figs. A.2, A.3, A.7, and A.9. Due to the long projection lever arm, the statistical precision on the intercept $\Delta\gamma_{\textsc{ess}}$ is good.
We observe in Fig. 6 that the ESS intercepts are close to zero but not always consistent with zero. Since these models shown in Fig. 6 contains no CME but only physics background, our results demonstrate that the $v_{2}$-induced background is unnecessarily zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$. Although the background may be suppressed, it is of no use because the amount of suppressed background is unknown, even whether it is over-suppressed or under-suppressed is unknown as shown in the lower panels of Fig. 6. It has been shown in [36] that CME signals would be preserved in the ESS intercept, however, such a conclusion or further studies of CME signal response in ESS is not helpful before the background issue is fully addressed. The reason that the $v_{2}$-induced backgrounds unnecessarily must be zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is because of the statistical fluctuation nature of the observables used in the ESS analysis. A model demonstration of this for the similar but simpler ES method (see Sect. II) can be found in Ref. [48]. There, the reason is clear, as elucidated in Sect. II and by the middle-panel cartoon of Fig. 1: the resonance (generally background contributing source) $v_{2}$ values are proportional to the overall $v_{2}$ of final-state particles, and when events are binned in the observed, statistically fluctuating $v_{2}^{\rm obs}$ of final-state particles, the average $\left\langle v_{2}\right\rangle$ values of the background sources are positive at $v_{2}^{\rm obs}=0$ [48]. The situation is much more complicated here in the ESS method where $\Delta\gamma$ is plotted against $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}$. As discussed in Sect. III.7, the $\hat{q}_{2,{\rm pair}}^{2}$ calculated by Eq. 9 is composed of several components: (i) a flow component from hydrodynamic particle pairs, which is related to the single particle flow in a non-trivial way [73], (ii) a flow component of the resonances (CME background contributing sources), (iii) a nonflow component between the two pairs used to calculate the two-pair cumulant by Eq. 9 that contain daughters from, for example, the same resonance decay, and (iv) a self-correlation component from the sharing of a common particle between the two pairs. Because of these complicated nature in the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ variable, it is unclear how the $v_{2,{\rm single}}$ and $\Delta\gamma$ would behave, individually and intertwiningly, as functions of $\hat{q}_{2,{\rm pair}}^{2}$. It would not be expected a priori that the $\Delta\gamma$ as a function of $v_{2,{\rm single}}$, binned in $\hat{q}_{2,{\rm pair}}^{2}$, would project to a zero intercept in the absence of CME.
It appears from the physics model studies shown in Fig. 6 (where CME is absent) that the value of $\Delta\gamma_{\textsc{ess}}$ can, indeed, be negative, zero, or positive, presumably dependent of model details (model descriptions can be found in Sects. III.1III.4). Note the results in Fig. 6 are averaged from 20–50% centralities. One obvious difference in the simulated events, even with the same model, is the event centrality or multiplicity. As discussed in Sect. III.7 and mentioned above, the $\hat{q}_{2,{\rm pair}}$ quantity contains a self-correlation component besides being sensitive to single-particle $v_{2}$, multi-particle cluster $v_{2}$, and nonflow. The self-correlations are caused by sharing of the same POI between two pairs of POIs, and thus its effect is inversely proportional to multiplicity. To gain insight, we show in Fig. 7 the ESS intercept in ampt and hydjet++ as a function of centrality in terms of the average POI multiplicity. The centralities are defined by the charged hadron multiplicity within $|\eta|<0.5$ to be 0–5%, 5–10%, 10–20%, $\cdots$ (10% size), 70–80%, and 80-100% of the total MB event sample. As seen from Fig. 7, the nonzero $\Delta\gamma_{\textsc{ess}}$ intercept becomes stronger in more peripheral collisions. Within the same model, the $\Delta\gamma_{\textsc{ess}}$ seems to follow the same trend in POI multiplicity. This explains why the hydjet++ $\Delta\gamma_{\textsc{ess}}$ in 20–50% centrality range is consistent with zero at 200 GeV and positive at 27 GeV. This may suggest that the root reason for the nonzero intercept $\Delta\gamma_{\textsc{ess}}$ is not collision energy or centrality, but POI multiplicity. The multiplicity dependence of the ESS intercept $\Delta\gamma_{\textsc{ess}}$ clearly shows that it is sensitive to background correlations. The different signs of $\Delta\gamma_{\textsc{ess}}$ between ampt (negative) and hydjet++ (positive) is presumably due to different physics implemented in these models.
Figure 7: ESS intercept multiplicity dependence. The ESS intercept $\Delta\gamma_{\textsc{ess}}$ as a function of the average POI multiplicity from peripheral 80–100% to central 0–5% Au+Au collisions at 200 and 27 GeV from ampt (upper panel) and hydjet++ (lower panel). The most peripheral point for AMPT 200 GeV is outside the frame.
To conclude this part of the discussion, our physics model results show that the ESS intercept is not an easily interpretable observable. Although the background in $\Delta\gamma_{\textsc{ess}}$ is strongly suppressed, one does not know how much background still remains in $\Delta\gamma_{\textsc{ess}}$, or even its sign. One does not know quantitatively what the ESS intercept exactly measures, so the ESS method is practically not useful to search for the CME.
IV.3 Toy Model Verification
One important difference among the models we studied is the various contributions from $K_{S}$. avfd and ampt keep the $K_{S}$ stable so there is no contamination from $K_{S}$ decay pions in the final state. epos4 and hydjet++, on the other hand, decay $K_{S}$ so there is maximum contamination in the final-state pions (as we have used all final-state pions of the model in our calculations). The intercepts from ampt at both 200 and 27 GeV are negative, and the intercept from avfd at 200 GeV is consistent with zero (avfd modeling at low energy is not available). The intercepts from epos4 at 200 and 27 GeV are both positive. The intercept from hydjet++ is consistent with zero at 200 GeV and positive at 27 GeV. In all models the strong-decay resonances (such as the $\rho^{0}$ resonance) are forced to decay. Relative particle abundances, for example, the $K_{S}$ over $\rho^{0}$ ratio, depends on energy; the lower the energy, the larger the low- to high-mass ratio. While very much dependent on models, it seems that $\Delta\gamma_{\textsc{ess}}$ increases with increasing $K_{S}/\rho^{0}$ ratio. It is thus conceivable, as resonances are a main background contribution to $\Delta\gamma$, that the relative resonance abundances could also be important for the ESS intercept.
To test the hypothesis that resonance mixtures may influence the robustness of the ESS method, we use toy model simulations, where we can easily alter the event content to examine the consequences. In real data analysis, a fraction of $K_{S}$ hadrons decay into pions that are reconstructed as primordial particles, often defined experimentally as those with a reconstructed distance of closest approach (dca) from the reconstructed primary collision point to be within a certain cut, typically a few centimeters. In the STAR experiment, such a fraction is on the order of 2% [49], and is used as default in Toy Model I (described in Sect. III.5). To alter the event content, we vary the accepted fraction of $K_{S}$ to 10%, 20%, and 30%.
In addition, we note in Fig. 2 that all the physics models contain a mass continuum of OS pairs in excess of SS pairs. We have therefore included a mass continuum in Toy Model II as described in Sect. III.6. The accepted $K_{S}$ fraction is kept at 2%. Adding the mass continuum is another way to alter the event content.
Figure A.14 shows the ESS results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum). It is interesting to notice that the dependence of $\Delta\gamma$ on $v_{2,{\rm single}}$ in those $\hat{q}_{2,{\rm pair}}^{2}$-binned events is not linear for Toy Model II. This suggests that the linear dependence hypothesis, which is essential for the linear extrapolation in ESS, may not be generally valid.
Figure A.14 shows the ESE results from Toy Model I (default 2% $K_{S}$), Toy Model I (10% $K_{S}$), and Toy Model II (mass continuum) with POIs from $0.3<|\eta|<1$ and $q_{2}^{2}$ from $|\eta|<0.3$. There seems to be an indication of nonlinear behavior in the Toy Model II result, similar to that in the ESS result in Fig. A.14 right panel.
Figure 8: Toy model ESS and ESE intercepts. Shown are intercepts (upper panel) and intercepts divided by the inclusive $\Delta\gamma$ value (lower panel) from the ESS (blue points) and ESE (red and pink points) methods using the toy models. Toy model inputs are from parameterizations to experimental data of 30–40% Au+Au collisions at 200 GeV. For ESS, the POI acceptance is $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs. For ESE, the POI acceptance is $0.3<|\eta|<1$, and the $\hat{q}_{2}^{2}$ is calculated from particles in $|\eta|<0.3$ (red points) or $3<|\eta|<4$ for three of the studied cases (pink points); The $p_{\rm T}$ range for both POIs and particles used for $\hat{q}_{2}^{2}$ calculations is $0.2<p_{\rm T}<2$ GeV/$c$. The rightmost red point is outside the frame.
Figure 8 summarizes the intercepts and the intercepts divided by the overall $\Delta\gamma$ values from the toy model studies. The intercepts from the ESS method (blue points) are all positive and appear to vary over the different event contents. This may confirm our hypothesis that the ESS method is sensitive to the details of the event makeup because of the complex intertwining of the $\Delta\gamma$, $v_{2,{\rm single}}$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}$.
The intercepts from the ESE method are shown in the red markers. The statistical precision is high because the toy models are computationally inexpensive. The ESE intercepts are consistent with zero for the two lower $K_{S}$-contamination cases studied in Toy Model I. However, the ESE intercepts from the two higher $K_{S}$-contamination cases of Toy Model I (positive) and the intercept from Toy Model II (negative) are inconsistent with zero. We postulate the cause to be nonflow correlations between POIs in $0.3<|\eta|<1$ and the $q_{2}^{2}$ calculated from $|\eta|<0.3$ in our toy model studies because resonance decay daughters can end up in both regions which are not far apart. (We have mentioned this nonflow effect in the $q_{2}^{2}$ event selection in our physics model discussion of Fig. 5, though the statistical uncertainties there are large.) In addition, the effect of adding a mass continuum is different from simply increasing the $K_{S}$ contamination, presumably because of the differing decay kinematics and opening angles.
To reduce this nonflow effect, we classify events in the ESE method using $q_{2}^{2}$ calculated from forward/backward pseudorapidity regions of $3<|\eta|<4$, far displaced from the POI’s region of $0.3<|\eta|<1$. The results are shown in Fig. A.14 for Toy Model I (20% $K_{S}$), Toy Model I (30% $K_{S}$), and Toy Model II (mass continuum). The ESE intercepts and the intercepts divided by the overall $\Delta\gamma$ values are shown in the pink markers in Fig. 8 with good statistical precision. These intercepts are consistent with zero, as one would expect in the absence of nonflow. These results highlight the importance to use $q_{2}^{2}$ calculated in pseudorapidity range far distanced from that of the POIs.
V summary
The charge-dependent azimuthal correlator $\Delta\gamma$ is commonly used to search for the chiral magnetic effect (CME), a parity and charge-parity violating phenomenon predicted by quantum chromodynamics. The $\Delta\gamma$ correlator is affected by a major physics background induced by elliptic flow anisotropy $v_{2}$ in relativistic heavy ion collisions. The event-shape engineering (ESE) and event-shape selection (ESS) methods have been exploited to search for the CME by projecting $\Delta\gamma$ to the measured $v_{2}$ of vanishing value, relying on dynamical fluctuations and statistical fluctuations of $v_{2}$, respectively.
We have conducted a systematic study using four physics models (avfd, ampt, epos4, and hydjet++) describing heavy ion collisions. It is found, with no CME signal, that the ESS intercept ($\Delta\gamma_{\textsc{ess}}$) can be negative, zero, or positive depending on the models and multiplicities of the events, and the ESE intercept ($\Delta\gamma_{\textsc{ese}}$) is mostly consistent with zero, as expected, albeit inherently large statistical uncertainties.
It is further found, using toy models, that $\Delta\gamma_{\textsc{ess}}$ appears to depend on the details of the event content, for example, the relative abundances of resonances and the presence of a mass continuum. Given that the event-shape selection variable $\hat{q}_{2,{\rm pair}}^{2}$, the single particle $v_{2,{\rm single}}$, and the three-point correlator $\Delta\gamma$ in $\hat{q}_{2,{\rm pair}}^{2}$-classified events all depend on the particles of interest (POIs), which are made of not only primordial particles but also all decay products of electromagnetic- and strong-decay resonances and partial products of weak-decay resonances, we conjecture that the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is a complicated function of all elements that make up the event. These elements include resonance abundances relative to primordial pions, resonance $v_{2}$’s and $p_{\rm T}$ spectra, as well as primordial particle $v_{2}$ and $p_{\rm T}$ spectra. Depending on the detailed physics of these elements, $\Delta\gamma_{\textsc{ess}}$ can be positive, negative, or zero. The issue is, unfortunately, that it is unknown, needless to say proven, what combinations of all those elements in the event would give the ideal $\Delta\gamma_{\textsc{ess}}=0$ in the absence of CME. This makes it impractical to search for CME using the ESS method.
The ESE method, on the other hand, has much more clarity: the $\Delta\gamma_{\textsc{ese}}$ intercept of flow-induced backgrounds must be zero in the absence of nonflow. However, the ESE method is not without pitfalls. There exist nonflow correlations between particles, and when those nonflow correlations are present between POIs and the $\hat{q}_{2}^{2}$ variable classifying events, then the ESE intercept can also be sensitive to non-CME physics. The issue is similar in nature to the ESS method–the relevant variables are now intertwined. One way to avoid/reduce such nonflow effects is to separate POIs and particles to calculate $\hat{q}_{2}^{2}$ well in pseudorapidity. Furthermore, the dynamical fluctuations in $v_{2}$ are typically not large, so a long extrapolation to the $v_{2}=0$ intercept is required in the ESE method. This results in a relatively large uncertainty in the $\Delta\gamma_{\textsc{ese}}$ intercept, and large statistics are required for a precision measurement by the ESE method. The statistics issue, however, can easily be resolved, at least conceptually, by accumulating more data.
Acknowledgment
This work is supported in part by the U.S. Department of Energy (Grant No. DE-SC0012910).
Appendix A Simulation Plots
This appendix compiles all simulation plots of $\Delta\gamma$ as functions of $v_{2}$ analyzed by the ESS and ESE methods from physics models as well as from toy models.
Figure A.1: avfd ESS results. Three values of $n_{5}/s=0$ (left panel), 0.1 (center panel), and 0.2 (right panel) are implemented in the avfd simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV, with approximately $2\times 10^{7}$ events each for the 30–40% centrality. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$.
Figure A.2: avfd ESE results. The same avfd data as in Fig. A.2 is used. The $\Delta\gamma$ is plotted as a function of $mean{v_{2}}$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
Figure A.3: ampt ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by ampt, with approximately $4.2\times 10^{8}$ and $1.8\times 10^{9}$ events for each centrality at each energy. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$
Figure A.4: ampt ESE results. The same ampt data as in Fig. A.3 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). The red line is a first-order polynomial fit to all data points.
Figure A.5: ampt ESE results with forward/backward $\hat{q}_{2}^{2}\{2\}$. As same as Fig. A.5 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in the forward/backward pseudorapidity region of $3<|\eta|<5$ for 200 GeV and $2<|\eta|<3$ for 27 GeV.
Figure A.6: epos4 ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by epos4, with approximately $1.6\times 10^{7}$ and $8.5\times 10^{6}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<2$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
Figure A.7: epos4 ESE results. The same epos4 data as in Fig. A.7 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<2$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
Figure A.8: hydjet++ ESS results. Shown are three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV (upper panels) and at 27 GeV (lower panels) simulated by hydjet++, with approximately $2.5\times 10^{8}$ and $5.7\times 10^{8}$ events for each centrality, respectively. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10).
Figure A.9: hydjet++ ESE results. The same hydjet++ data as in Fig. A.9 is used. The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10).
Figure A.10: hydjet++ ESE results (three-particle correlator). As same as Fig. A.9 except that the three-particle correlator method is used to calculate $\Delta\gamma$ (Eq. 13) and the two-particle cumulant is used to calculate $\left\langle v_{2}\right\rangle$ (Eq. 8), instead of using the known impact parameter direction in these calculations.
Figure A.11: hydjet++ ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). As same as Fig. A.9 except that the event selection variable $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6) is computed from particles in forward/backward pseudorapidity regions of $3<|\eta|<5$ for 200 GeV and $2<|\eta|<3$ for 27 GeV.
Figure A.12: Toy model ESS results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel) and of 10% (middle panel), and Toy Model II with mass continuum (right panel). The $K_{S}$, $\rho^{0}$, and primordial pion inputs are from parameterizations to experimental data of 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The numbers of events are $2\times 10^{8}$ for each panel. The $\Delta\gamma$ is plotted as a function of $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ (Eqs. 5,9). POIs are from acceptance $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$, and the event selection variable $\hat{q}_{2,{\rm pair}}^{2}\{2\}$ is computed from the same POIs. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $v_{2,{\rm single}}$ (Eq. 10). The red line is a first-order polynomial fit in the range of $0<v_{2,{\rm single}}<0.2$.
Figure A.13: Toy model ESE results. Toy Model I with the accepted $K_{S}$ fraction of the default 2% (left panel, $2.8\times 10^{9}$ events) and of 10% (middle panel, $2.0\times 10^{9}$ events), and Toy Model II with mass continuum (right panel, $2.2\times 10^{8}$ events). The $\Delta\gamma$ is plotted as a function of $\left\langle v_{2}\right\rangle$ in events binned in $\hat{q}_{2}^{2}\{2\}$ (Eqs. 5,6). POIs are from acceptance $0.3<|\eta|<1$, and the event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $|\eta|<0.3$, both with $0.2<p_{\rm T}<2$ GeV/$c$. The model’s known impact parameter direction $\psi=0$ is taken as the EP in calculating $\Delta\gamma$ (Eqs. 2,3) and $\left\langle v_{2}\right\rangle$ (Eq. 10). Other information is as same as those in Fig. A.14. The red line is a first-order polynomial fit to all data points.
Figure A.14: Toy model ESE results (forward/backward rapidity $\hat{q}_{2}^{2}$). Toy Model I with the accepted $K_{S}$ fraction of 20% (left panel, $2.2\times 10^{8}$ events) and of 30% (middle panel, $2.2\times 10^{8}$ events), and Toy Model II with mass continuum (right panel, $10^{8}$ events). The event selection variable $\hat{q}_{2}^{2}\{2\}$ is computed from particles in $3<|\eta|<4$. Other information is as same as those in Fig. A.14.